# Exploring Kolmogorov-Arnold networks for realistic image sharpness assessment

### v2

@@ Abstract
Score prediction is crucial in realistic image sharpness assessment after informative features are collected. Recently, Kolmogorov-Arnold networks (KANs) have been developed and witnessed remarkable success in data fitting. This study presents Taylor series based KAN (TaylorKAN). Then, different KANs are explored on four realistic image databases (BID2011, CID2013, CLIVE, and KonIQ-10k) for score prediction by using 15 mid-level features and 2048 high-level features. When setting support vector regression as the baseline, experimental results indicate KANs are generally better or competitive, TaylorKAN is the best on three databases using mid-level feature input, while KANs are inferior on CLIVE when high-level features are used. This is the first study that explores KANs for image quality assessment. It sheds lights on how to select and improve KANs on related tasks.
@@
Kolmogorov-Arnold network, image sharpness assessment, image quality, machine learning.
@@ I Introduction
Blind image sharpness/blurriness assessment (BISA) plays an important role in media quality assurance. It could automate real-time processing of large volumes of images without any reference information, allow for dynamic adjustment of image sharpness and visual fidelity in video streaming, and enhance the quality of user experience [1]. Specifically, BISA can guide the image restoration process by offering a metric to optimize algorithms for sharpening and deblurring [2].
@@
Score prediction is indispensable when informative features have been prepared for image sharpness presentation. Support vector regression (SVR) and multi-layer perceptron (MLP) are preferred. Li $et$ $al$ craft multi-scale sharpness-aware features, and SVR performs score rating [3]. Yu $et$ $al$ design a convolutional neural network (CNN) for blurriness estimation [4], and SVR and MLP are compared [5]. Tang $et$ $al$ quantify image sharpness from intrinsic structures, and random forest is used for feature pooling and quality rating [6]. Liu $et$ $al$ craft many statistical features, and a three-hidden-layered MLP scores image quality [7]. Liu $et$ $al$ design orientation-aware features for SVR-based score prediction [8]. Chen $et$ $al$ weight local binary pattern features in spatial domain and entropy and gradient features in spectral domain, and SVR predicts the quality scores from perception features [9]. Yu $et$ $al$ construct mid-level features, and MLP, SVR, general regression neural network and elastic net are evaluated [10].
@@
For score prediction, a CNN can be treated as an image-based feature extractor followed by a MLP in an optimization procedure. Many CNNs have been proposed. Zhu $et$ $al$ retrieve prior knowledge shared among distortions, and the network is then fine-tuned for quality prediction [11]. Song $et$ $al$ integrate domain knowledge of human visual system and natural scene statistics into a deep network for scoring [12]. Li $et$ $al$ utilize a pre-trained network to quantify the discrepancy and image quality [13]. Huang $et$ $al$ investigate the inherent relationship between the attributes and the category via graph convolution network for semantic attribute reasoning and image quality estimation [14]. Zhang $et$ $al$ optimize a deep CNN on multiple databases by a hinge constraint on learning uncertainty [15]. Li and Huo consider global and local features in multi-scale distortion perception and introduce the feedback mechanism [16]. Chen $et$ $al$ design multi-scale spatial pyramid pooling that combines block attention and image understanding for improved accuracy and generalization [17]. Sun $et$ $al$ extract low-level visual features and high-level semantic information and a staircase structure is designed for hierarchical feature integration and quality-aware representation [18]. Zhao $et$ $al$ enable learning representation by a pretext task in a self-supervised learning manner, and a quality-aware contrastive loss is employed to learn distortions from image patches [19]. Zhang $et$ $al$ develop a multitask learning scheme for image quality assessment, image scene classification and distortion type identification, and the model parameter sharing and the loss weighting are determined automatically [20]. Wu $et$ $al$ fuse multi-stage semantic features for image quality prediction, and the features are rectified using a novel multi-level channel attention module [21].
@@
Inspired by the Kolmogorov-Arnold theorem (KAT), a novel module called Kolmogorov-Arnold Networks (KAN) has been proposed [22]. Except for the activation functions as MLP on nodes, KAN appends learnable activation functions on edges between the nodes of successive layers. Improved representation capacity has been witnessed in function approximation, data fitting, knowledge representation and collaborative filtering [23, 24, 25]. Later, different KAN variants are implemented using different mathematical functions as the activation functions on edges, such as FourierKAN-GCF [25], BSRBF_KAN [26], ChebyKAN [27], and WavKAN [28].
@@
Despite remarkable success in knowledge representation and data fitting, little is known about KANs for score prediction in BISA. To bridge this gap, this study explores different KANs for image blurriness estimation. Firstly, Taylor series based KAN (TaylorKAN) is proposed. Then, mid-level features and deeply learned features of four realistic databases are prepared. After that, MLP, SVR and six KANs are evaluated. Results indicate that KANs are generally better than SVR and MLP, and TaylerKAN is the best on three databases when mid-level features are used.
@@ II-B KAN variants
The flexibility of KAN structure allows various implementations using different activation functions on edges. The proposed TaylorKAN and other involved variants are described.
@@ II-B1 TaylorKAN
Taylor series (Eq. 4) can represent a function $f(x)$ as an infinite sum of terms from the values of its derivatives $f^{(n)}$ at a point $a$. TaylorKAN is based on Taylor series, and the coefficients are learned during model training.
@@ II-B3 ChebyKAN
ChebyKAN employs Chebyshev polynomials for function approximation [27]. It attempts to minimize the maximum error in polynomial approximation and enables ChebyKAN suited for high accuracy and stability. Eq. 6 shows Chebyshev polynomials defined by the recurrence relation.
@@ II-B4 HermiteKAN
Hermite polynomials are ideal for approximating these Gaussian-like functions due to the recurrence relations and orthogonality. HermiteKAN is based on Hermite polynomials (Eq. 7).
@@ II-B5 JacobiKAN
JacobiKAN uses Jacobi polynomials. The polynomials are orthogonal to the weight function $(1-x)^{\alpha}(1+x)^{\beta}$ and flexible in handling diverse boundary conditions.
@@ II-B6 WavKAN
Wavelets are adept at capturing local variations in functions and provide flexibility and adaptability in complex pattern modeling. The general wavelet transformation form is shown as below,
@@
where $a$ and $b$ correspond to the scaling and translation parameters. Since different wavelet basis are available, in this study, the Mexican Hat wavelets are used in WavKAN,
@@ II-B7 Other variants
Other KANs are available [24]. However, limited by computing resource, above-mentioned light-weight KANs are evaluated.
@@ III-B2 High-level features
Using pre-trained ResNet50 as an extractor $f$, it generates a $d$-dimensional feature vector ($d=2048$) to an input image $I_{i}$. Thereby, deeply learned high-level feature matrix $N$ is derived as shown in Eq. 12
@@ III-C Score prediction
MLP ($mlp$) is with one hidden layer of 10 nodes. Its parameters are optimized by minimizing the difference between the model $R_{mlp}$ output and the ground truth $Y$.
@@ III-D Performance criterion
The performance is evaluated using Pearson linear correlation coefficient (PLCC), Spearman rank order correlation coefficient (SRCC) and Kendalls rank correlation coefficient (KRCC). Higher values indicate better performance.
Notably, PLCC is computed after a five-parameter nonlinear mapping between objective and subjective scores. The mapping function is shown as below,
$f(s)=q_{1}(\frac{1}{2}-\frac{1}{1+e^{q_{2}(s-q_{3})}})+q_{4}s+q_{5},$ (15)
where $s$ denotes the predicted score, $f(s)$ is the mapped score, and $\{q_{i}\}_{i=1}^{5}$ are the fitting parameters. The PLCC value is then calculated between subjective scores and mapped scores.
@@ III-E Implementation details
A database is randomly splitted into three subsets for model training (70% samples), validation (15% samples) and testing (15% samples). All the score prediction models are run on the same data sets.
TaylorKAN is embedded with 2 hidden layers ([15, 8, 4, 1]) for mid-level feature input and with 3 hidden layers ([2048, 512, 256, 128, 1]) for high-level feature input. Specifically, quadratic approximation is used in TaylorKAN. Other models are used with default parameter settings. The project code is available online (www.github.com/CUC-Chen/kanBISA).
@@ IV Results and Discussion
Score prediction results using mid-level features and high-level features are shown in Table II and Table III, respectively. Table IV compares the current study with several state-of-the-art achievement. The highest metric values are in boldface.
@@ IV-A BISA using mid-level features
Table II shows that compared to SVR, all the KAN models are better on BID2011, TaylorKAN and HermiteKAN are superior on CID2013, TaylorKAN, BSRBF_KAN, JacobiKAN, and HermiteKAN achieve slightly higher PLCC on CLIVE, and the KANs lead to competitive results on KonIQ-10k.
TABLE II: No-reference assessment results by using mil-level features
BID2011 CID2013 CLIVE KonIQ-10k
PLCC SRCC KRCC PLCC SRCC KRCC PLCC SRCC KRCC PLCC SRCC KRCC
SVR 0.564 0.561 0.391 0.831 0.822 0.683 0.589 0.561 0.392 0.764 0.732 0.535
MLP 0.589 0.588 0.418 0.839 0.835 0.681 0.572 0.542 0.381 0.752 0.721 0.526
BSRBF_KAN 0.697 0.705 0.519 0.833 0.802 0.600 0.601 0.514 0.360 0.742 0.667 0.481
ChebyKAN 0.711 0.769 0.560 0.753 0.792 0.606 0.522 0.454 0.306 0.752 0.675 0.488
HermiteKAN 0.721 0.779 0.579 0.815 0.822 0.634 0.581 0.522 0.368 0.747 0.678 0.490
JacobiKAN 0.719 0.778 0.567 0.792 0.830 0.631 0.572 0.475 0.324 0.749 0.676 0.489
WavKAN 0.655 0.685 0.486 0.850 0.824 0.641 0.606 0.498 0.343 0.762 0.690 0.502
TaylorKAN 0.757 0.750 0.550 0.859 0.843 0.666 0.608 0.488 0.342 0.719 0.640 0.458
It is observed that when using 15 mid-level features, the proposed TaylorKAN leads to the highest PLCC on BID2011, CID2013 and CLIVE, and slightly lower values on KonIQ-10k to the baseline. The reasons are manifold. Firstly, mid-level features intrinsically represent image sharpness. And thereby, all KANs, particularly TaylorKAN, obtain much higher values on BID2011. Secondly, the characteristic of image sharpness representation constrains the prediction performance of KANs on CID2013, CLIVE and KonIQ-10k, since these databases contains diverse distortions. It becomes understandable that KANs obtain inferior results on the largest, realistic database KonIQ-10k. Thirdly, the representation capacity of the mid-level features might be insufficient that poses difficulties on accurate score prediction.
IV-B BISA using deeply learned features
Setting SVR as the baseline model, Table III shows KANs achieve higher values on BID2011 and much lower values on CID2013, BSRBF_KAN and WavKAN obtain slight increase on PLCC on CLIVE, and KANs achieve superior performance on KonIQ-10k.
TABLE III: No-reference assessment results by using deeply learned features
BID2011 CID2013 CLIVE KonIQ-10k
PLCC SRCC KRCC PLCC SRCC KRCC PLCC SRCC KRCC PLCC SRCC KRCC
SVR 0.774 0.752 0.562 0.924 0.909 0.745 0.733 0.724 0.533 0.803 0.783 0.585
MLP 0.746 0.725 0.538 0.925 0.909 0.748 0.465 0.440 0.310 0.607 0.576 0.408
BSRBF_KAN 0.818 0.813 0.607 0.858 0.861 0.689 0.737 0.663 0.487 0.840 0.807 0.615
ChebyKAN 0.809 0.816 0.621 0.525 0.527 0.382 0.683 0.545 0.394 0.838 0.804 0.611
HermiteKAN 0.825 0.831 0.635 0.636 0.635 0.472 0.741 0.704 0.520 0.844 0.810 0.617
JacobiKAN 0.826 0.834 0.633 0.683 0.685 0.513 0.747 0.694 0.515 0.842 0.809 0.616
WavKAN 0.797 0.811 0.608 0.839 0.850 0.663 0.761 0.687 0.507 0.848 0.817 0.625
TaylorKAN 0.792 0.788 0.587 0.750 0.756 0.565 0.667 0.563 0.405 0.831 0.801 0.610
When using 2048 high-level features for image quality representation, TaylorKAN obtains slightly higher metric values on BID2011 and KonIQ-10k, and lower values on CID2013 and CLIVE to SVR. The inferior performance suggests that TaylorKAN may not be well-suited for handling high-dimensional features. Compared to the metric values derived from mid-level features, all KANs, except BSRBF_KAN, show significant decrease on CID2013, and their improvements on the other databases are not as substantial as those of SVR. Another reason might come from the representation capacity of quadratic approximation in TaylorKAN. Higher-order Taylor expansions may improve prediction accuracy, while simpler approximations can simplify complex problems, enhance the efficiency of the optimization process, and reduce the risk of over-fitting in score prediction.

### v3

@@ Abstract
Score prediction is crucial in evaluating realistic image sharpness based on collected informative features. Recently, Kolmogorov-Arnold networks (KANs) have been developed and witnessed remarkable success in data fitting. This study introduces the Taylor series-based KAN (TaylorKAN). Then, different KANs are explored in four realistic image databases (BID2011, CID2013, CLIVE, and KonIQ-10k) to predict the scores by using 15 mid-level features and 2048 high-level features. Compared to support vector regression, results show that KANs are generally competitive or superior, and TaylorKAN is the best one when mid-level features are used. This is the first study to investigate KANs on image quality assessment that sheds some light on how to select and further improve KANs in related tasks.
@@
Kolmogorov-Arnold network, TaylorKAN, image sharpness assessment, image quality, machine learning.
@@ I Introduction
Blind image sharpness/blurriness assessment (BISA) is crucial in media quality assurance. It enables real-time processing without reference images and supports dynamic adjustments to image sharpness and visual fidelity in video streaming, thereby enhancing the quality of the user experience. [1]. Specifically, BISA can be used to guide the restoration process by offering a metric to optimize algorithms for image sharpening [2].
@@
Score prediction is an indispensable step when informative features have been prepared for image sharpness representation. Support vector regression (SVR) and multi-layer perceptron (MLP) are preferred. Li $et$ $al$ craft multi-scale sharpness-aware features, and SVR performs score rating [3]. Yu $et$ $al$ design a convolutional neural network (CNN) for blurriness estimation [4], and SVR and MLP are used to improve the prediction performance [5]. Liu $et$ $al$ design orientation-aware features for SVR-based score prediction [6]. Chen $et$ $al$ weight the local binary pattern features in spatial domain and entropy and gradient features in spectral domain, and SVR predicts the quality scores from perception features [7]. Yu $et$ $al$ construct mid-level features, and MLP and SVR are evaluated [8].
@@
For score prediction, a CNN model can be generally treated as an image-based feature extractor followed by a MLP in end-to-end optimization. Zhu $et$ $al$. retrieve prior knowledge shared among distortions, and a deep net is fine-tuned for quality scoring [9]. Huang $et$ $al$. investigate the inherent relationship between the attributes and the categories via graph convolution network for attribute reasoning and quality estimation [10]. Zhang $et$ $al$. optimize a CNN on multiple databases by a hinge constraint on learning uncertainty [11]. Li and Huo consider multi-scale visual features and introduce the feedback mechanism [12]. Chen $et$ $al$. develop multi-scale spatial pooling and combine both block attention and image understanding for improved generalization [13]. Sun $et$ $al$. extract low-level features and high-level semantics, and a staircase structure is designed for hierarchical feature integration and quality-aware embedding [14]. Zhao $et$ $al$. enable representation learning via a pre-text self-supervised task and use a quality-aware contrastive loss to learn distortions [15]. Zhang $et$ $al$. design multi-task learning for quality assessment, scene classification and distortion identification [16]. Wu $et$ $al$. fuse multi-stage semantic features for no-reference image quality prediction, and before score rating, the features are rectified using multi-level channel attention [17].
@@
Inspired by the Kolmogorov-Arnold theorem (KAT), a novel module called Kolmogorov-Arnold Networks (KAN) has been proposed [18]. Except for the activation functions as MLP on nodes, KAN appends learnable activation functions on edges between the nodes of successive layers. Improved capacity has been shown in data fitting and knowledge representation [19]. Later, KAN variants are designed using different mathematical functions as the activation functions on edges [20].
@@
Despite remarkable success in knowledge representation and data fitting, little is known about KANs for score prediction in BISA. This study attempts to bridge this gap. Firstly, the Taylor series-based KAN (TaylorKAN) is introduced. Then, mid-level features and high-level features of four realistic databases are prepared. After that, MLP, SVR and 6 KANs are evaluated. Experimental results indicate that KANs are generally better than SVR and MLP, and TaylerKAN is the best when using mid-level features as image quality representation.
@@ II-B KAN variants
The flexibility of KAN structure allows for diverse implementations using different activation functions on edges. The TaylorKAN and other involved KANs are described.
@@ II-B1 TaylorKAN
Taylor series (Eq. 4) represent $f(x)$ as an infinite sum of terms from the values of its derivatives $f^{(n)}$ at a point $a$. In TaylorKAN, the coefficients are learned during model training.
@@ II-B3 ChebyKAN
ChebyKAN employs the Chebyshev polynomials for function approximation [23]. It attempts to minimize the maximum error in polynomial approximation for high accuracy and stability. Eq. 6 shows Chebyshev polynomials defined by the recurrence relation.
@@ II-B4 HermiteKAN
Hermite polynomials are ideal for approximating these Gaussian-like functions due to their recurrence relations and orthogonality. HermiteKAN is based on Hermite polynomials (Eq. 7).
@@ II-B5 JacobiKAN
JacobiKAN uses Jacobi polynomials [24]. The polynomials are orthogonal to the weight function and flexible in handling diverse boundary conditions.
@@ II-B6 WavKAN
Wavelets are adept at capturing local variations in functions, and they provide flexibility and adaptability in complex pattern modeling. The general wavelet transformation form is shown as below,
@@
where $a$ and $b$ correspond to the scaling and the translation parameters. Since different wavelet basis are available, in this study, the Mexican Hat wavelets are used in WavKAN [25],
@@ II-B7 Other variants
Other KANs are accessible [20]. However, limited by computing resources, above-mentioned light-weight KANs are evaluated in the current study.
@@ III-B2 High-level features
Using pre-trained ResNet50 [30] as an extractor $f$, it generates a $d$-dimensional feature vector ($d=2048$) to an input image $I_{i}$. Thereby, deeply learned high-level feature matrix $N$ is derived as shown in Eq. 12.
@@ III-C Score prediction
MLP ($mlp$) is designed with different numbers of the hidden layers in terms of different feature inputs. Its parameters are optimized by minimizing the difference between the output of model $R_{mlp}$ and the ground truth $Y$.
@@ III-D Performance criterion
The performance is evaluated by using Pearson linear correlation coefficient (PLCC) and Spearman rank order correlation coefficient (SRCC). Higher values indicate better performance.
Specifically, PLCC is computed after a five-parameter nonlinear mapping between objective and subjective scores (Eq. 15) in which $s$ denotes the predicted score, $f(s)$ is the mapped score, and $\{q_{i}\}_{i=1}^{5}$ are the fitting parameters. PLCC values are calculated between subjective scores and mapped scores.
$f(s)=q_{1}(\frac{1}{2}-\frac{1}{1+e^{q_{2}(s-q_{3})}})+q_{4}s+q_{5}$ (15)
Training time is used to evaluate the efficiency of involved score prediction models. The training time for SVR is recorded in seconds ($s$), and the other models are measured in iterations per second ($i/s$), where a lower value indicates poorer computational efficiency.
@@ III-E Implementation details
In each experiment, a database is randomly split into three subsets for training (70% samples), validation (15% samples) and testing (the remaining samples) of BIQA indicators.
For fair comparison, the KANs and MLP are configured with 3 hidden layers ([15, 26, 18, 12, 1]) for mid-level inputs and 4 hidden layers ([2048, 1536, 1024, 256, 128, 1]) for high-level feature inputs. TaylorKAN is implemented with quadratic approximation, and the other KAN models use default settings. To SVR, the radial basis function (RBF) kernel is used, and the other parameters are set with default values.
During training, an early stopping mechanism is used even though 500 iterations are pre-defined. The patience parameter is set to 20, and thus, the training will be terminated early if the validation loss fails to improve for 20 consecutive epochs. Throughout the process, the metrics to the learning rate that yields the highest sum of PLCC and SRCC are reported.
The codes are implemented on an Ubuntu operating system (version 22.04) using Python (version 3.12), PyTorch (version 2.3.0), and CUDA (version 12.1). The algorithms are executed on a GPU (A100-PCIE-40GB) with 72 GB RAM. The project is available at https://github.com/CUC-Chen/KAN4IQA.
@@ IV Results and Discussion
Experimental results using mid-level features and high-level features are respectively shown in Table II and III where the denotes the training time of SVR in $s$. Table IV compares the current study with several state-of-the-art achievement. In the tables, the best metric values are boldfaced.
@@ IV-A BISA using mid-level features
Table II shows that compared to SVR, TaylorKAN achieves better prediction on BID2011, CID2013 and KonIQ-10k, and its PLCC value is higher on CLIVE. The other KAN models obtain better performance on BID2011, competitive results on CID2013 and KonIQ-10k, but worse prediction on CLIVE. Among the KANs, ChebyKAN is the fastest, achieving over 30 $i/s$ on BID2011 and CID2013, and TaylorKAN demonstrates slightly lower but comparable computational efficiency.
TABLE II: BISA by using 15 mil-level features
BID2011 [26] CID2013 [27] CLIVE [28] KonIQ-10k [29]
PLCC SRCC time ($i/s$) PLCC SRCC time ($i/s$) PLCC SRCC time ($i/s$) PLCC SRCC time ($i/s$)
SVR 0.619 0.617 0.049 0.834 0.810 0.024 0.630 0.592 0.117 0.746 0.691 12.39
MLP 0.744 0.729 33.66 0.808 0.791 39.62 0.649 0.552 18.06 0.753 0.682 2.082
BSRBF_KAN [22] 0.675 0.680 33.23 0.845 0.795 12.99 0.562 0.479 6.347 0.725 0.650 1.463
ChebyKAN [23] 0.700 0.703 34.86 0.808 0.826 37.73 0.570 0.447 18.36 0.749 0.680 2.224
HermiteKAN 0.651 0.740 19.29 0.825 0.845 20.99 0.566 0.502 9.655 0.754 0.671 1.118
JacobiKAN [24] 0.709 0.789 16.29 0.808 0.775 20.12 0.545 0.519 9.546 0.753 0.689 1.074
WavKAN [25] 0.715 0.730 22.74 0.827 0.827 26.60 0.559 0.482 13.23 0.759 0.685 1.448
TaylorKAN (ours) 0.756 0.782 28.44 0.871 0.851 34.45 0.668 0.582 15.32 0.766 0.699 1.927
When using the 15 mid-level features for quality representation, TaylorKAN achieves the best performance, attributed to several factors. Firstly, it surpasses SVR by leveraging the Taylor series for hierarchical approximation, and meanwhile, a multi-layer network is implemented for deeper feature learning. Secondly, it outperforms other involved KANs, possibly because quadratic approximation enables more precise feature fitting. Thirdly, it achieves higher metric values than MLP, since the former uses additional activation functions on edges. At last, the simplicity of the quadratic approximation enables TaylorKAN to perform data fitting at a relatively faster pace.
IV-B BISA using high-level features
Setting SVR as the baseline, Table III suggests that KANs achieve higher values on BID2011, worse results on CID2013 and CLIVE, and close performance on KonIQ-10k in general. Notably, BSRBF_KAN and WavKAN are comparable to SVR on CID2013 and CLIVE, while TaylorKAN, BSRBF_KAN, JacobiKAN and HermiteKAN are superior on KonIQ-10k. It is found that BSRBF_KAN, WavKAN and TaylorKAN run at less than 10 $i/s$ on BID2011 and CID2013.
TABLE III: BISA by using 2048 high-level features
BID2011 [26] CID2013 [27] CLIVE [28] KonIQ-10k [29]
PLCC SRCC time ($i/s$) PLCC SRCC time ($i/s$) PLCC SRCC time ($i/s$) PLCC SRCC time ($i/s$)
SVR 0.786 0.782 0.355 0.860 0.882 0.447 0.751 0.712 1.492 0.839 0.800 119.22
MLP 0.750 0.780 23.46 0.796 0.825 18.32 0.637 0.554 12.77 0.808 0.763 1.038
BSRBF_KAN [22] 0.811 0.795 6.296 0.828 0.820 6.247 0.733 0.649 3.768 0.841 0.809 0.435
ChebyKAN [23] 0.821 0.812 20.79 0.630 0.665 18.61 0.662 0.587 11.41 0.824 0.790 1.368
HermiteKAN 0.822 0.814 12.53 0.604 0.687 10.45 0.670 0.647 6.804 0.839 0.804 0.803
JacobiKAN [24] 0.820 0.806 11.38 0.596 0.605 10.65 0.733 0.651 6.426 0.842 0.803 0.699
WavKAN [25] 0.767 0.735 1.178 0.844 0.856 0.994 0.752 0.676 0.587 0.810 0.777 0.067
TaylorKAN (ours) 0.797 0.813 3.285 0.788 0.780 2.796 0.696 0.598 1.644 0.850 0.811 0.188
When applying the 2048 high-level features for image quality representation, TaylorKAN obtains slightly higher metric values on BID2011 and KonIQ-10k, while worse results on CID2013 and CLIVE than SVR. The sub-optimal performance indicates that it remains challenging for the KANs to handle high-dimensional features. On the one hand, the KAT (Eq. 1) implies that the representation of a high-dimensional function requires a large number of uni-variate functions. Consequently, the accumulation of errors and over-fitting may cause unstable data-driven approximation. On the other hand, whether these high-level features primarily learned for object recognition can be directly utilized to represent image quality has not yet been determined. Therefore, proper post-processing strategies, such as fine-tuning [9] and rectification [17], become important for smoothly transferring these deeply learned features into the representation space of image quality.
Comparison of Table II and Table III reveals that the 2048 high-level features provide more effective quality representation than the 15 mid-level features for score prediction, except on the CID2013 database. Most KANs, such as BSRBF_KAN, ChebyKAN and JacobiKAN, achieve better performance when high-level features are used, while the utilization of the 2048 features dramatically increases the computational time. On the one hand, it should be acknowledged the mid-level features are less effective than the high-level features due to their limited quantity and capacity for quality representation. Specifically, the mid-level features are primarily designed for scoring image sharpness [8], while the high-level features are hierarchically and deeply learned for general recognition task [30]. On the other hand, the KANs perform generally better on CID2013 when using the mid-level features. For example, in the case of JacobiKAN, using mid-level features leads to PLCC 0.808 and SRCC 0.775, whereas using high-level features yields PLCC 0.596 and SRCC 0.605. This finding indicates that the mid-level features are more suitable for the CID2013 database.